# P.O.Box 19395-5531, Tehran, Iran

### v1

@@ 4 Summary
In the first part of this work, we propose a black hole metric with a static horizon embedded in FLRW cosmology. This metric has the following properties:
@@
$e)$ If we fix the sign of $g_{tr}$ to be negative, the resulting metric satisfies the null convergence condition outside the horizon but violates it inside the horizon.
@@
Since our proposed metric violates NEC inside the horizon, we drop one of our assumptions and see that it is in principle possible to find a cosmological metric with a static horizon that satisfies NEC everywhere.
In the second part of this work we considered a more realistic framework with a constant equation of state (for the cases of matter and radiation) and show that we have no black hole solutions for the case of $\omega\neq 0$. For the case of matter with $\omega=0$, we derived the black hole metric consistent with this assumption. The crucial feature of this metric is that despite having an asymptotically vanishing Hubble parameter, we have no singularities at the horizon.
We also studied the case of matter$+$cosmological constant where we assumed the pressure component remains equal to the constant cosmological value. In all of these cases we show that the black hole horizon can asymptotically (for large $t$) tend to a static horizon.
@@ Acknowledgment
We thank Mohammad Mehdi Sheikh-Jabbari for comments and discussions.

### v2

@@ 4 Summary
We find that it is technically much easier to find white hole metrics in an expanding cosmology or a black hole metric in a contracting cosmology. This is due to the fact that the sign of $h$ in (18) is negative for a black hole and positive for an expanding cosmology. Therefore, to find a black hole metric embedded in expanding cosmology, we need to have a sign change in $h$ in the vicinity of the black hole relative to the asymptotic cosmology.
Due to this complexity, we first studied the more trivial case of a white hole in expanding cosmology and then dealt with the issue of the sign change in $h$ by generalizing the form of the metric. The generalization is effectively a curvature-like behavior (relative to the flat FLRW metric).
In the first part of this work, we propose white and black hole metrics with static horizons embedded in FLRW cosmology. These metrics have the following properties:
@@
$e)$ It violates NEC. But we show that it is in principal possible to drop assumption c and find metrics with static horizons that satisfy NEC as well.
@@
In the second part of this work we considered a more realistic framework with a constant equation of state (for the cases of matter and radiation) and first study the white hole case due to its simplicity. We show that we have no non-trivial solutions for the case of $\omega\neq 0$. For the case of matter with $\omega=0$, we derived the white hole metric consistent with this assumption. The crucial feature of this metric is that, despite having an asymptotically vanishing Hubble parameter, we have no singularities at the horizon.
We also studied the case of matter plus cosmological constant, where we assumed the pressure component remains equal to the constant cosmological value. In all of these cases, we show that the white hole horizon can asymptotically (for large $t$) tend to a static horizon.
Finally we considered the case of black hole immersed in cosmological matter (with and without cosmological constant) with a velocity function $h$ that changes sign in the vicinity of the black hole.
@@ Acknowledgment
We thank Mohammad Mehdi Sheikh-Jabbari for comments and discussions. We specially thank Grigorii E. Volovik for mentioning his important work that helped us improve this manuscript.